# Workshop on Mechanistic Interpretability

### v4

@@ Appendix E Experimental Details
We run all experiments using H100 GPUs, and we use the sentence tokeniser from the nltk library [1] to preprocess all datasets. For all RAG models, i.e., Qwen2-1.5B-Instruct, Qwen2-7B-Instruct [39], Gemma2-2B-Instruct and Gemma2-9B-Instruct [30], we use their standard chat templates to construct the prompt, i.e., using the context and query as a user’s message.
@@
Appendix G Compute-accuracy Trade-off Between Different Baselines and Our ARC-JSD
We mainly compare the compute-accuracy trade-off between different baselines and our proposed ARC-JSD when attributing responses to relevant context. As Figure 5 and 6 show, our ARC-JSD method can achieve up to 3-fold speedup compared to other baselines. In addition, ARC-JSD is consistently Pareto-optimal over different LLM backbone sizes.
@@ Appendix G Compute-accuracy Trade-off Between Different Baselines and Our ARC-JSD
Appendix H Examples of ARC-JSD Context Attribution
@@
Example of ARC-JSD attributions for Qwen2-7B-IT on MuSiQue Query What is the date of birth of the person who was part of Ratata? Response (attributed statement is highlighted) The date of birth for Mauro Scocco, who was part of the pop group Ratata, is given as September 11, 1962. attribution
Top-1 Source (from the context)
@@
His mother’s name was not recorded in history. Title: Mauro Scocco Mauro Scocco (born 11 September 1962) is a Swedish pop artist of Italian descent. He has been described as “one of the sharpest songwriters in Sweden”.
Example of ARC-JSD attributions for Gemma2-2B-IT on TyDi QA Query When was the friends song was released which is taken from the fourth studio album of canadian singer and song writer Justin Bieber? Response (attributed statement is highlighted) The song "Friends" was released on **August 17, 2017**. It’s important to note that this is from Justin Bieber’s fourth studio album, "Purpose". attribution
Top-1 Source (from the context)
It was written by Bieber, Julia Michaels, Justin Tranter and BloodPop, with production handled by BloodPop. The song was released through GENPOP Corporation, RBMG Records, School Boy Records, Def Jam Recordings and Republic Records on August 17, 2017. It is Bieber’s first single credited as lead artist to be released in over a year, the previous being "Company", which was released in March 2016 for his studio album "Purpose" (2015).
@@
Example of ARC-JSD attributions for Gemma2-9B-IT on Hotpot QA Query Tom Thumb Food & Pharmacy is part of the southern division of a grocery company based in what city? Response (attributed statement is highlighted) The answer is **Boise, Idaho**. The text states that Tom Thumb Food & Pharmacy is part of the Southern division of Albertsons, and Albertsons Companies LLC is based in Boise, Idaho. attribution
Top-1 Source (from the context)
@@
It is also known in Brazil as "Chapolin", "Vermelhinho" ("Little Red One") and "Polegar Vermelho" ("Red Thumb") in allusion to the famous fairy tale character Tom Thumb. Title: Albertsons Albertsons Companies LLC is an American grocery company founded and based in Boise, Idaho. It is privately owned and operated by investors, including Cerberus Capital Management.
@@
Appendix I JSD-based Mechanistic Insights for Located Attention Heads and MLPs
@@
Appendix J Comparisons of JSD with KL, Wasserstein, TV and MMD in Detail
@@
Most existing baselines, e.g., ContextCite [6], SelfCite [3] and AttriBoT [19], use direct log-probability or KL divergence as metric for context attribution. However, these metrics drop diverges if the masked run assigns $\approx 0$ probability to the token, which is sensitive to highly-skewed token frequencies. Moreover, if JSD is replaced with KL in the Eq. 1 and Eq. 2, it will bring some influence to the attribution impact:
@@
    • Semantic blindness: For the example: “Paris” -> “Lyon” (same POS, both cities) and “Paris” -> “Party”. At the token level, any substitution has unit cost, so replacing “Paris” with either “Lyon” or “Party” is equally cheap, despite radically different semantic consequences. With subword tokenisation, the cost becomes tokeniser-dependent. Character-level edit distance differentiates orthography (e.g., “Party” is closer to “Paris” than “Lyon”), which misaligns with factual attribution
@@
Appendix K Case Studies of Attention and MLP’s Contribution for Each Response Token

### v5

@@ Appendix E Experimental Details
We run all experiments using H100 GPUs, and we use the sentence tokeniser from the nltk library Bird et al. (2009) to preprocess all datasets. For all RAG models, i.e., Qwen2-1.5B-Instruct, Qwen2-7B-Instruct Yang et al. (2024), Gemma2-2B-Instruct and Gemma2-9B-Instruct Team et al. (2024), we use their standard chat templates to construct the prompt, i.e., using the context and query as a user’s message.
@@
Appendix G The Use of Large Language Models (LLMs) Statement
This work was finished and written without the help of LLMs.
Appendix H Compute-accuracy Trade-off Between Different Baselines and Our ARC-JSD
We mainly compare the compute-accuracy trade-off between different baselines and our proposed ARC-JSD when attributing responses to relevant context. As Figure 8 and 9 show, our ARC-JSD method can achieve up to 3-fold speedup compared to other baselines. In addition, ARC-JSD is consistently Pareto-optimal over different LLM backbone sizes.
@@ Appendix H Compute-accuracy Trade-off Between Different Baselines and Our ARC-JSD
Appendix I Examples of ARC-JSD Context Attribution
@@
Example of ARC-JSD attributions for Qwen2-7B-IT on MuSiQue Query What is the date of birth of the person who was part of Ratata? Response (attributed statement is highlighted) The date of birth for Mauro Scocco, who was part of the pop group Ratata, is given as September 11, 1962. Gold Answer 11 September 1962 Attribution
Top-1 Source (JSD: 18.64)
@@
Mauro Scocco (born 11 September 1962) is a Swedish pop artist of Italian descent.
Top-2 Source (JSD: 0.24)
Title: Ratata (band)
Top-3 Source (JSD: 0.11)
Ratata was a Swedish pop group fronted by Swedish vocalist Mauro Scocco formed in 1980.
Top-4 Source (JSD: 0.10)
Title: Mauro Scocco
Top-5 Source (JSD: 0.03)
After Ratata, Scocco has continued as a solo artist since.
Top-6 Source (JSD: 0.02)
He has been described as ”one of the sharpest songwriters in Sweden”.
Example of ARC-JSD attributions for Gemma2-2B-IT on TyDi QA Query Who founded the Burntisland Shipbuilding Company? Response (attributed statement is highlighted) Amos and Wilfrid Ayre. Gold Answer Amos and Wilfrid Ayre Attribution
Top-1 Source (JSD: 41.86)
Brothers Amos and Wilfrid Ayre founded Burntisland Shipbuilding Co. in 1918 as a First World War emergency shipyard.
Top-2 Source (JSD: 0.12)
Its yard at Burntisland West Dock had four berths and capacity to build ships up to 450 feet (140m) long[1] and up to 59 feet (18m) beam.
@@
Example of ARC-JSD attributions for Gemma2-9B-IT on Hotpot QA Query Tom Thumb Food & Pharmacy is part of the southern division of a grocery company based in what city? Response (attributed statement is highlighted) The answer is **Boise, Idaho**. The text states that Tom Thumb Food & Pharmacy is part of the Southern division of Albertsons, and Albertsons Companies LLC is based in Boise, Idaho. Gold Answer Boise Attribution
Top-1 Source (JSD: 6.73)
@@
It is also known in Brazil as ”Chapolin”, ”Vermelhinho” (”Little Red One”) and ”Polegar Vermelho” (”Red Thumb”) in allusion to the famous fairy tale character Tom Thumb. Title: Albertsons Albertsons Companies LLC is an American grocery company founded and based in Boise, Idaho. It is privately owned and operated by investors, including Cerberus Capital Management.
Top-2 Source (JSD: 6.14)
It operate under the names Tom Thumb traditional grocery stores; Flagship Tom Thumb high end stores, usually in affluent areas. It makes up part of the Southern division of Albertsons. It is (as of May 2015) the number two supermarket in the competitive…
Appendix J ARC-JSD used on medical and legal tasks
We further test our method by randomly selecting examples from the medical or legal datasets, e.g., PubMedQA, MedQuAD, and LegalBench, using the general instruction-tuning LLM backbones.
Based on those randomly selected examples below, we can find that our ARC-JSD is still applicable to the medical and legal domains. The top-1 located context with the highest JSD score correctly supports model’s response. In addition, we can further understand how model’s thinking about other context segments by looking at their JSD scores in the ranking list.
Medical Example 1: Disease Symptoms and Characteristics Context Parkinson’s disease (PD) is a progressive neurodegenerative disorder that primarily affects movement. The hallmark motor symptoms include bradykinesia (slowness of movement), resting tremor, and rigidity. In addition to motor symptoms, patients often experience a wide range of non-motor symptoms, such as cognitive impairment, sleep disturbances, and autonomic dysfunction. The underlying pathology involves the loss of dopaminergic neurons in the substantia nigra, a region of the midbrain. Query What are the primary motor symptoms associated with Parkinson’s disease? Response (attributed statement is highlighted) The primary motor symptoms associated with Parkinson’s disease (PD) include:
1. Bradykinesia - Slowness of movement.
2. Resting tremor - A persistent shaking or trembling sensation during rest.
3. Rigidity - Difficulty moving certain muscles, especially when lying down or sleeping.
These symptoms can significantly impact daily activities and quality of life for individuals with PD.
Attribution
Top-1 Source (JSD: 4.89)
The hallmark motor symptoms include bradykinesia (slowness of movement), resting tremor, and rigidity.
Top-2 Source (JSD: 0.63)
Parkinson’s disease (PD) is a progressive neurodegenerative disorder that primarily affects movement.
Top-3 Source (JSD: 0.15)
In addition to motor symptoms, patients often experience a wide range of non-motor symptoms, such as cognitive impairment, sleep disturbances, and autonomic dysfunction.
Top-4 Source (JSD: 0.11)
The underlying pathology involves the loss of dopaminergic neurons in the substantia nigra, a region of the midbrain.
Medical Example 2: Adverse Effects and Contraindications Context ACE inhibitors (e.g., lisinopril, ramipril) are commonly prescribed for hypertension and heart failure. A well-known side effect is a persistent, dry cough, which is thought to be mediated by bradykinin accumulation. More serious, though less common, adverse effects include hyperkalemia (high potassium levels) and angioedema (rapid swelling). These drugs are strictly contraindicated in pregnancy due to their teratogenic effects, particularly during the second and third trimesters. Query Why should ACE inhibitors not be used during pregnancy? Response (attributed statement is highlighted) ACE inhibitors, such as lisinopril and ramipril, are commonly prescribed for hypertension and heart failure. However, they have been found to cause several serious adverse effects when taken during pregnancy, including hyperkalemia (high potassium levels) and angioedema (rapid swelling). Additionally, these drugs are strictly contraindicated in pregnancy due to their teratogenic effects, particularly during the second and third trimesters. Therefore, it is important to avoid using ACE inhibitors during pregnancy to minimize the risk of harm to both mother and baby. Attribution
Top-1 Source (JSD: 20.31)
These drugs are strictly contraindicated in pregnancy due to their teratogenic effects, particularly during the second and third trimesters.
Top-2 Source (JSD: 16.71)
More serious, though less common, adverse effects include hyperkalemia (high potassium levels) and angioedema (rapid swelling).
Top-3 Source (JSD: 9.83)
ACE inhibitors (e.g., lisinopril, ramipril) are commonly prescribed for hypertension and heart failure.
Top-4 Source (JSD: 1.05)
A well-known side effect is a persistent, dry cough, which is thought to be mediated by bradykinin accumulation.
Legal Example 1: Criminal Procedure (Miranda Rights) Context The Miranda v. Arizona ruling by the Supreme Court established that custodial interrogation is inherently coercive. To safeguard the Fifth Amendment privilege against self-incrimination, law enforcement officials must inform a suspect in custody of their rights before questioning begins. These rights include the right to remain silent, the warning that anything said can be used against them in court, the right to an attorney, and the right to have an attorney appointed if they cannot afford one. A waiver of these rights must be made voluntarily, knowingly, and intelligently. Query What specific rights must a suspect in custody be informed of before being questioned, according to the Miranda ruling? Response (attributed statement is highlighted) According to the Miranda ruling, a suspect in custody must be informed of their rights before being questioned, including the following:
1. The right to remain silent.
2. The warning that anything said can be used against them in court.
3. The right to an attorney.
4. The right to have an attorney appointed if they cannot afford one.
These rights must be informed voluntarily, knowingly, and intelligently.
Attribution
Top-1 Source (JSD: 8.65)
These rights include the right to remain silent, the warning that anything said can be used against them in court, the right to an attorney, and the right to have an attorney appointed if they cannot afford one.
Top-2 Source (JSD: 6.10)
A waiver of these rights must be made voluntarily, knowingly, and intelligently.
Top-3 Source (JSD: 0.73)
To safeguard the Fifth Amendment privilege against self-incrimination, law enforcement officials must inform a suspect in custody of their rights before questioning begins.
Top-4 Source (JSD: 0.54)
The Miranda v. Arizona ruling by the Supreme Court established that custodial interrogation is inherently coercive.
Legal Example 2: Intellectual Property (Copyright Fair Use) Context Section 107 of the U.S. Copyright Act provides for the ”fair use” of a copyrighted work without the owner’s permission for purposes such as criticism, comment, news reporting, teaching, scholarship, or research. To determine if a particular use is ”fair,” courts evaluate four statutory factors. These factors are: (1) the purpose and character of the use, including whether it is commercial or for non-profit educational purposes; (2) the nature of the copyrighted work; (3) the amount and substantiality of the portion used; and (4) the effect of the use upon the potential market for the work. Query What are the four factors a court uses to determine if the use of copyrighted material qualifies as ’fair use’? Response (attributed statement is highlighted) The four factors that a court uses to determine if the use of copyrighted material qualifies as ’fair use’ are:
1. The purpose and character of the use, including whether it is commercial or for non-profit educational purposes.
2. The nature of the copyrighted work.
3. The amount and substantiality of the portion used.
4. The effect of the use upon the potential market for the work.
These factors help courts assess whether the use of copyrighted material falls within the bounds of fair use and does not infringe on the copyright holder’s rights.
Attribution
Top-1 Source (JSD: 21.45)
These factors are: (1) the purpose and character of the use, including whether it is commercial or for non-profit educational purposes; (2) the nature of the copyrighted work; (3) the amount and substantiality of the portion used; and (4) the effect of the use upon the potential market for the work.
Top-2 Source (JSD: 0.41)
To determine if a particular use is ”fair,” courts evaluate four statutory factors.
Top-3 Source (JSD: 0.27)
Section 107 of the U.S. Copyright Act provides for the ”fair use” of a copyrighted work without the owner’s permission for purposes such as criticism, comment, news reporting, teaching, scholarship, or research.
@@
Appendix K JSD-based Mechanistic Insights for Located Attention Heads and MLPs
@@
Appendix L Comparisons of JSD with KL, Wasserstein, TV and MMD in Detail
@@
Most existing baselines, e.g., ContextCite Cohen-Wang et al. (2024), SelfCite Chuang et al. (2025) and AttriBoT Liu et al. (2024), use direct log-probability or KL divergence as metric for context attribution. However, these metrics drop diverges if the masked run assigns $\approx 0$ probability to the token, which is sensitive to highly-skewed token frequencies. Moreover, if JSD is replaced with KL in the Eq. 1 and Eq. 2, it will bring some influence to the attribution impact:
@@
    • Semantic blindness: For the example: “Paris” $\rightarrow$ “Lyon” (same POS, both cities) and “Paris” $\rightarrow$ “Party”. At the token level, any substitution has unit cost, so replacing “Paris” with either “Lyon” or “Party” is equally cheap, despite radically different semantic consequences. With subword tokenisation, the cost becomes tokeniser-dependent. Character-level edit distance differentiates orthography (e.g., “Party” is closer to “Paris” than “Lyon”), which misaligns with factual attribution
@@
Appendix M Case Studies of Attention and MLP’s Contribution for Each Response Token